OpenStudy (anonymous):

The focus for this parabola is (-1,0). x=(1/4)y^2 A. True B. False

OpenStudy (dumbcow):

didn't you already do this one...

OpenStudy (dumbcow):

the last post i said...the focus width = 1/(4a) = 1/4*1/4 = 1

OpenStudy (dumbcow):

another way to check is the focus is always inside the parabola. So if it opens up/right, the focus is above or right of vertex. if it opens down/left, the focus is below or left of vertex

OpenStudy (anonymous):

so this is true?

OpenStudy (dumbcow):

this open to the right, therefore the focus must be to the right of vertex. Is (-1,0) to the right of (0,0) ?

OpenStudy (anonymous):

no

OpenStudy (dumbcow):

good, that means (-1,0) is not the focus from my earlier post you should already know that the focus of x=1/4y^2 is (1,0)